Suppose you are working with a data set that is normally distributed with a mean of 400 and a standard deviation of 20. Determine the value of x such that 60% of the values are greater than x.
Answer by stanbon(75887)

Suppose you are working with a data set that is normally distributed with a mean of 400 and a standard deviation of 20. Determine the value of x such that 60% of the values are greater than x.
-----
Find the z-value with a left tail of 0.40
invNorm(0.40) = -0.2533
----
Find the corresponding x-value using x = z*s + u
x = -0.2533*20 + 400 = 394.93
==================================
Cheers,
Stan H.
=================